OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 8/x^2 + 4

OpenStudy (vishweshshrimali5):

There can be several such functions. I am giving you one example here and then try finding out one yourself: Let f(x) = 8x + 4 and g(x) = \(\cfrac{1}{x^2}\) Then, f(g(x)) = 8\(\cfrac{1}{x^2}\) + 4 = y

OpenStudy (anonymous):

I don't understand

OpenStudy (vishweshshrimali5):

Okay let me help you.

OpenStudy (vishweshshrimali5):

First of all you have two options you can either assume f(x) or g(x).

OpenStudy (vishweshshrimali5):

For example, in the example I wrote above , I assumed f(x) first, f(x) = 8x + 4 What I actually did was I observed y carefully. See if I forget the 1/x^2 part in y and replace it with "?" I get; y = 8? + 4 Right ?

OpenStudy (vishweshshrimali5):

@Andshedances1234 ??

OpenStudy (anonymous):

I'm here, im trying to think lol. Why would you forget the x^2?

OpenStudy (vishweshshrimali5):

So that I can easily assume f(x) and g(x)

OpenStudy (vishweshshrimali5):

I am just replacing 1/x^2 by some other symbol "?"

OpenStudy (anonymous):

Okay, so we assumed f(x)? What about g(x)

OpenStudy (vishweshshrimali5):

You can easily obtain g(x) once you have assumed f(x). See: I have y = 8? + 4 Ok?

OpenStudy (anonymous):

right

OpenStudy (vishweshshrimali5):

Good now for my example, I assumed f(x) = "?" Okay ?

OpenStudy (anonymous):

ok

OpenStudy (vishweshshrimali5):

Now, what we actually do by doing f(g(x)) is in place of x in f(x) we put g(x). Do you get this ? for example, take a very different example, suppose f(x) = 2x Now, if I replace x by g(x), then f(x) will change to f(g(x)) So, f(g(x)) = 2g(x)

OpenStudy (vishweshshrimali5):

Do you get this part ?